May I have some help with the following equation?
Find all real solutions of the equation.
x^8 + 15x^4=16

Answer by checkley71(8403)

X^8+15X^4-16=0
(X^4+16)(X^4-1)=0
(X^4+16)(X^2+1)(X^2-1)=0
(X^4+16)(X^2+1)(X+1)(X-1)=0